# No Language is an Island: Unifying Chinese and English in Financial Large Language Models, Instruction Data, and Benchmarks

### v1

@@ A Examples of Context and Answer in Financial Task
BigData22
Context: The recent trend of \$intc is as follows. date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2020-09-14, -1.1, 0.6, -1.1, 0.3, 0.3, -0.4, 1.3, 1.0, 0.5, 0.1, -0.2; 2020-09-15, -0.4, 1.2, -0.5, 1.2, 1.2, -1.1, -0.1, -0.1, -0.6, -1.0, -1.3; 2020-09-16, 0.3, 1.3, -0.4, 0.7, 0.7, -1.5, -0.9, -0.7, -1.2, -1.5, -1.9; 2020-09-17, -1.9, 0.3, -2.0, -0.1, -0.1, -0.9, -1.2, -0.5, -0.9, -1.3, -1.7; 2020-09-18, 0.9, 1.2, -1.7, -0.9, -0.9, 0.2, -0.4, 0.4, 0.1,-0.4,-0.7; 2020-09-21, -0.7, 0.0, -1.8, -0.3, -0.3, 0.7, -0.1, 0.7, 0.5, 0.0, -0.3; 2020-09-22, -0.1, 0.5, -1.0, 0.5, 0.5, 0.2, -0.4, 0.1, 0.1, -0.4, -0.7; 2020-09-23, 2.1, 2.7, -0.3, -2.3, -2.3, 1.9, 1.7, 2.1, 2.3, 2.0, 1.6;2020-09-24, -1.3, 1.0, -1.5, 0.7, 0.7, 0.7, 1.1, 1.0, 1.6, 1.3, 0.9; 2020-09-25, -2.0, 0.7, -2.4, 1.6, 1.6, -0.8, -0.4, -0.6, 0.0, -0.2, -0.5.
@@
NER
Context: The proceeds of the equipment advances will be used solely to reimburse Borrower for the purchase of eligible equipment.
@@
FINER-ORD
Context: McPhail , 50 , is angry and perplexed by politicians clamouring to sign a free trade deal with China that allows the importation of Chinese workers .
@@
ECTSUM
Context: And I’m confident that our disciplined approach to operating the business will result in our continued success throughout the balance of fiscal 2021. Coupled with the cost reduction actions we recently implemented, …, Our Test business delivered a really solid Q1 by significantly beating our internal expectations and delivering an EBITDA margin of nearly 13% versus 11% last Q1.
@@
EDTSUM
Context: ORLANDO, Fla., June 4, 2020 PRNewswire–Darden Restaurants, Inc., (NYSE: DRI) plans to release its fiscal 2020 fourth quarter financial results…For more information, please visit www.darden.com. Logo: https:mma.prnewswire.com media 340562 darden-restaurants-inc-logo.jpg SOURCE Darden Restaurants, Inc.: Financial Related Links http:www.Darden.com
@@
German
Context: The client has attributes: X1: A14, X2: 12, X3: A32, X4: A40, X5: 1386, X6: A63, X7: A73, X8: 2, X9: A92, X10: A101, X11: 2, X12: A122, X13: 26, X14: A143, X15: A152, X16: 1, X17: A173, X18: 1, X19: A191, X20: A201.
@@
Australian
Context: The client has attributes: A1: 0.0, A2: 68.67, A3: 15.0, A4: 2.0, A5: 10.0, A6: 9.0, A7: 0.0, A8: 1.0, A9: 1.0, A10: 14.0, A11: 0.0, A12: 2.0, A13: 0.0, A14: 3377.0.
@@
FOMC
Context: Nonetheless, given a recovery in U. S. domestic demand approximating their current forecasts, growth in imports likely would exceed that of exports by a wide margin over the forecast horizon.
@@ C Example of Chinese Dataset Translation
D Human Evaluation for Human-written Prompts
Table 7-8 presents the example of human evaluation for domain-expert written prompts on various financial specific task
@@ D Human Evaluation for Human-written Prompts
Sentiment Analysis (FinSA)
确定所提供的金融新闻文章中的句子的情绪,识别其中的情绪是积极,消极,还是中性的。只需回答’积极’、’中性’或’消极’。
涨空间异常巨大,中铁装配,实在是未来的大牛股啊
积极
积极
3 3 3 3
Sentiment Matching (FinSM)
你需要判断给出的两个金融文本表达的意思是否相似,请回答是或否。
花呗咋扫不上/花呗扫一扫不能付钱
0 1 1 0.7
News Classification (FinNC)
请根据金融报道的内容,判断其涉及范围属于’中国’、’国际’还是’外国’?
伯恩斯坦:首予康菲石油(COP.N)跑赢大盘评级,目标价为98美元。
外国
国际
1 3 1 1.7
Negative Judgment (FinNJ)
根据所给的金融新闻及实体,你需要判断所给实体是否含有负面的消息,你只需简单回答’有’或者’无’。
自从号称750亿交易量的唐小僧倒下…几乎每天都有几家网贷公司宣布清盘、提现困难、经侦介入等。实体:唐小僧;多米
3 3 3 3
Answer Selection (FinAS)
根据所提出的金融问题,请从以下四个选项中选择最合适的一个。你的输出应该是:’A’、’B’、’C’ 或’D’。
如果一个竞争性市场位于长期均衡状态中,那么所有的厂商()。 A,采用完全相同的生产工艺; B,具有同一最低平均的成本; C,都能获得经济利润; D,以上全对
D
D
3 3 2 2.7
Stock Prediction (FinSP)
请根据新闻对股票数据的影响判断该公司的股票运动走势是跑赢大盘、中性还是表现不佳。
高新兴公告,实控人拟减持不超过6%。日期 开盘价 闭盘价,2020-09-23, 5.56, 5.52;2020-09-24, 5.47, 5.22; … ; 2020-09-29, 5.02, 5.03
表现不佳
无法预测
0 1 2 1
Headline Classification (FinHC)
下面所给内容是否提到了黄金商品的上涨趋势?请回答是或否。
金价在印度的主要城市上涨
3 3 2 2.7
Question Answering (FinQA)
你需要分析金融文本,根据内容回答相关问题。
【亿邦动力讯】10月21日消息,洛丁森完成B轮融资,…,下文中发生的事件类型有哪些?
企业融资
无法确定
0 0 0 0
Event Detection (FinED)
阅读以下金融领域的公告,判断所有的事件类型及其对应主体。请以:’事件类型, 事件主体’的格式回复。其中事件类型应该在[’信批违规’,…,’涉嫌欺诈’]中。
劲嘉股份(002191)因虚假陈述遭证监会处罚浩云科技(300448)股东披露减持计划
信批违规, 劲嘉股份
信批违规, 劲嘉股份
3 3 3 3
Entity Recognition (FinER)
在下面所给金融语句中,识别指明个人(’PER’)、组织(’ORG’)或地点(’LOC’)的特定命名实体。
报告期内,本公司除中国银保监会批准的经营范围内的金融担保业务外,没有其他需要披露的重大担保事项。
中国银保监会, ORG
中国银保监会, ORG;中国, LOC
2 2 3 2.3
Relationship Extraction (FinRE)
请仔细分析所给金融报道和实体对,然后在[’合并’, ’竞争’,…]中选择能准确描述该实体对关系的选项。
张裕与卡斯特集团合作的牵线人陈光因嫌分红太少而有意退出。实体对:卡斯特集团-张裕
合作
合作关系受损
1 2 3 2
Text Summarization (FinTS)
你的任务是对所给的金融文本进行一句话摘要,要求尽量简短。
前段时间,佳能宣布关闭经营了32年的珠海工厂,引发网络关注。…而优厚的补偿,也是佳能自愿选择的“体面离开”。(新京报)
佳能撤离珠海,不宜说成“恶意赔偿”
佳能关闭珠海工厂,提供优厚补偿,引发争议。
1 2 3 2
@@
Determine the mood of the sentences in the financial news article provided, identifying whether the mood in them is positive, negative, or neutral. Simply answer ’positive’, ’neutral’, or ’negative’.
Up space is exceptionally huge, China Railway assembly, is really the future of the big bull stock ah
positive
positive
3 3 3 3
Sentiment Matching (FinSM)
@@
You need to decide whether the two financial texts given express similar meanings, answer yes or no.
Why can’t I pay/ I can’t pay by scanning.
yes
no
0 1 1 0.7
News Classification (FinNC)
@@
Based on the content of the financial report, please determine whether its scope is ’Chinese’, ’international’ or ’foreign’?
Bernstein: first Outperform rating on ConocoPhillips (COP.N), \$98 price target.
foreign
international
1 3 1 1.7
Negative Judgment (FinNJ)
Based on the given financial news and entity, you need to determine whether the given entity contains negative news or not, you simply need to answer ’yes’ or ’no’.
@@
Since the fall of Tang Xiaosheng, which claimed a transaction volume of 75 billion … Almost every day, there are several online lending companies announced liquidation, cash withdrawal difficulties, the investigation intervention, etc. Entity: Tang Xiaosheng; Domi
yes
yes
3 3 3 3
Answer Selection (FinAS)
Based on the financial question posed, choose the most appropriate of the following four options. Your output should be: ’A’, ’B’, ’C’ or ’D’.
@@
If a competitive market is in long-run equilibrium, then all firms () A, use exactly the same production process; B, have the same minimum average cost; C, are economically profitable; D, all of the above.
D
D
3 3 2 2.7
Stock Prediction (FinSP)
Determine whether the company’s stock movement trend is outperforming, neutral, or underperforming based on the impact of the news on the stock data.
@@
Gaoxin announced that the real controller intends to reduce its holdings by no more than 6%. Date Opening price Closing price : 2020-09-23, 5.56, 5.52 ; 2020-09-24, 5.47, 5.22; … ; 2020-09-29, 5.02, 5.03
underperformance
unpredictable
0 1 2 1
Headline Classification (FinHC)
@@
Does the content given below refer to an upward trend in gold commodities? Please answer yes or no.
Gold prices rise in India’s major cities
yes
yes
3 3 2 2.7
Question Answering (FinQA)
You will need to analyze financial texts and answer relevant questions based on the content.
@@
[YPC] October 21, 2011 - Lodgingsen closed its Series B financing… What are the types of events that occur in the following article?
corporate financing
indeterminate
0 0 0 0
Event Detection (FinED)
@@
Read the following announcements in the financial sector and determine all event types and their corresponding subjects. Please respond in the format :’Event Type, Event Subject’. Where the event type should be in [’credit approval violation’, …’ Suspected Fraud’].
Jinjia shares (002191) due to misrepresentation by the Securities and Futures Commission penalized haoyun science and technology (300448) shareholders disclosure of shareholding reduction plan
Letter approval violation, Jinjia shares
Letter approval violation, Jinjia shares
3 3 3 3
Entity Recognition (FinER)
In the financial statement given below, identify the specific named entity that specifies a person (’PER’), organization (’ORG’), or location (’LOC’).。
@@
During the reporting period, the Company did not have any other material guarantee matters requiring disclosure, except for the financial guarantee business within the scope of operation approved by the China Banking and Insurance Regulatory Commission.
China Banking and Insurance Regulatory Commission, ORG
China Banking Regulatory Commission, ORG; China, LOC
2 2 3 2.3
Relationship Extraction (FinRE)
@@
Carefully analyze the given financial reports and entity pairs, and then among [’Merger’, ’Competition’, …] Choose the option that accurately describes the relationship of the entity pair.
Chen Guang, the matchmaker of Zhangyu’s partnership with Custer Group, intends to quit because he thinks the dividend is too small. Entity Pair: Custer Group - Zhang Yu
collaborate
Impaired partnership
1 2 3 2
Text Summarization (FinTS)
Your task is to make a one-sentence summary of the given financial text, keeping it as short as possible.
@@
Some time ago, Canon announced the closure of its Zhuhai plant, which had been in operation for 32 years, triggering concern on the Internet…. The generous compensation is also Canon’s voluntary choice to "leave with dignity". (New Beijing News)
Canon’s withdrawal from Zhuhai should not be labeled as "malicious compensation".
Canon closes Zhuhai plant, offers generous compensation, sparks controversy.
1 2 3 2

### v2

@@ A Examples of Context and Answer in Financial Task
BigData22 Context: The recent trend of \$intc is as follows. date, open, high, low, close, adj-close, inc-5, inc-10, inc-15, inc-20, inc-25, inc-30; 2020-09-14, -1.1, 0.6, -1.1, 0.3, 0.3, -0.4, 1.3, 1.0, 0.5, 0.1, -0.2; 2020-09-15, -0.4, 1.2, -0.5, 1.2, 1.2, -1.1, -0.1, -0.1, -0.6, -1.0, -1.3; 2020-09-16, 0.3, 1.3, -0.4, 0.7, 0.7, -1.5, -0.9, -0.7, -1.2, -1.5, -1.9; 2020-09-17, -1.9, 0.3, -2.0, -0.1, -0.1, -0.9, -1.2, -0.5, -0.9, -1.3, -1.7; 2020-09-18, 0.9, 1.2, -1.7, -0.9, -0.9, 0.2, -0.4, 0.4, 0.1,-0.4,-0.7; 2020-09-21, -0.7, 0.0, -1.8, -0.3, -0.3, 0.7, -0.1, 0.7, 0.5, 0.0, -0.3; 2020-09-22, -0.1, 0.5, -1.0, 0.5, 0.5, 0.2, -0.4, 0.1, 0.1, -0.4, -0.7; 2020-09-23, 2.1, 2.7, -0.3, -2.3, -2.3, 1.9, 1.7, 2.1, 2.3, 2.0, 1.6;2020-09-24, -1.3, 1.0, -1.5, 0.7, 0.7, 0.7, 1.1, 1.0, 1.6, 1.3, 0.9; 2020-09-25, -2.0, 0.7, -2.4, 1.6, 1.6, -0.8, -0.4, -0.6, 0.0, -0.2, -0.5.
@@
NER Context: The proceeds of the equipment advances will be used solely to reimburse Borrower for the purchase of eligible equipment.
@@
FINER-ORD Context: McPhail , 50 , is angry and perplexed by politicians clamouring to sign a free trade deal with China that allows the importation of Chinese workers .
@@
ECTSUM Context: And I’m confident that our disciplined approach to operating the business will result in our continued success throughout the balance of fiscal 2021. Coupled with the cost reduction actions we recently implemented, …, Our Test business delivered a really solid Q1 by significantly beating our internal expectations and delivering an EBITDA margin of nearly 13% versus 11% last Q1.
@@
EDTSUM Context: ORLANDO, Fla., June 4, 2020 PRNewswire–Darden Restaurants, Inc., (NYSE: DRI) plans to release its fiscal 2020 fourth quarter financial results…For more information, please visit www.darden.com. Logo: https:mma.prnewswire.com media 340562 darden-restaurants-inc-logo.jpg SOURCE Darden Restaurants, Inc.: Financial Related Links http:www.Darden.com
@@
German Context: The client has attributes: X1: A14, X2: 12, X3: A32, X4: A40, X5: 1386, X6: A63, X7: A73, X8: 2, X9: A92, X10: A101, X11: 2, X12: A122, X13: 26, X14: A143, X15: A152, X16: 1, X17: A173, X18: 1, X19: A191, X20: A201.
@@
Australian Context: The client has attributes: A1: 0.0, A2: 68.67, A3: 15.0, A4: 2.0, A5: 10.0, A6: 9.0, A7: 0.0, A8: 1.0, A9: 1.0, A10: 14.0, A11: 0.0, A12: 2.0, A13: 0.0, A14: 3377.0.
@@
FOMC Context: Nonetheless, given a recovery in U. S. domestic demand approximating their current forecasts, growth in imports likely would exceed that of exports by a wide margin over the forecast horizon.
@@ C Example of Chinese Dataset Translation
D Human Evaluation for Expert-written Prompts
Table 8-9 presents the example of human evaluation for domain-expert written prompts on various financial specific task
@@ D Human Evaluation for Expert-written Prompts
Sentiment Analysis (FinSA) 确定所提供的金融新闻文章中的句子的情绪,识别其中的情绪是积极,消极,还是中性的。只需回答’积极’、’中性’或’消极’。 涨空间异常巨大,中铁装配,实在是未来的大牛股啊 积极 积极 3 3 3 3
Sentiment Matching (FinSM) 你需要判断给出的两个金融文本表达的意思是否相似,请回答是或否。 花呗咋扫不上/花呗扫一扫不能付钱 0 1 1 0.7
News Classification (FinNC) 请根据金融报道的内容,判断其涉及范围属于’中国’、’国际’还是’外国’? 伯恩斯坦:首予康菲石油(COP.N)跑赢大盘评级,目标价为98美元。 外国 国际 1 3 1 1.7
Negative Judgment (FinNJ) 根据所给的金融新闻及实体,你需要判断所给实体是否含有负面的消息,你只需简单回答’有’或者’无’。 自从号称750亿交易量的唐小僧倒下…几乎每天都有几家网贷公司宣布清盘、提现困难、经侦介入等。实体:唐小僧;多米 3 3 3 3
Answer Selection (FinAS) 根据所提出的金融问题,请从以下四个选项中选择最合适的一个。你的输出应该是:’A’、’B’、’C’ 或’D’。 如果一个竞争性市场位于长期均衡状态中,那么所有的厂商()。 A,采用完全相同的生产工艺; B,具有同一最低平均的成本; C,都能获得经济利润; D,以上全对 D D 3 3 2 2.7
Stock Prediction (FinSP) 请根据新闻对股票数据的影响判断该公司的股票运动走势是跑赢大盘、中性还是表现不佳。 高新兴公告,实控人拟减持不超过6%。日期 开盘价 闭盘价,2020-09-23, 5.56, 5.52;2020-09-24, 5.47, 5.22; … ; 2020-09-29, 5.02, 5.03 表现不佳 无法预测 0 1 2 1
Headline Classification (FinHC) 下面所给内容是否提到了黄金商品的上涨趋势?请回答是或否。 金价在印度的主要城市上涨 3 3 2 2.7
Question Answering (FinQA) 你需要分析金融文本,根据内容回答相关问题。 【亿邦动力讯】10月21日消息,洛丁森完成B轮融资,…,下文中发生的事件类型有哪些? 企业融资 无法确定 0 0 0 0
Event Detection (FinED) 阅读以下金融领域的公告,判断所有的事件类型及其对应主体。请以:’事件类型, 事件主体’的格式回复。其中事件类型应该在[’信批违规’,…,’涉嫌欺诈’]中。 劲嘉股份(002191)因虚假陈述遭证监会处罚浩云科技(300448)股东披露减持计划 信批违规, 劲嘉股份 信批违规, 劲嘉股份 3 3 3 3
Entity Recognition (FinER) 在下面所给金融语句中,识别指明个人(’PER’)、组织(’ORG’)或地点(’LOC’)的特定命名实体。 报告期内,本公司除中国银保监会批准的经营范围内的金融担保业务外,没有其他需要披露的重大担保事项。 中国银保监会, ORG 中国银保监会, ORG;中国, LOC 2 2 3 2.3
Relationship Extraction (FinRE) 请仔细分析所给金融报道和实体对,然后在[’合并’, ’竞争’,…]中选择能准确描述该实体对关系的选项。 张裕与卡斯特集团合作的牵线人陈光因嫌分红太少而有意退出。实体对:卡斯特集团-张裕 合作 合作关系受损 1 2 3 2
Text Summarization (FinTS) 你的任务是对所给的金融文本进行一句话摘要,要求尽量简短。 前段时间,佳能宣布关闭经营了32年的珠海工厂,引发网络关注。…而优厚的补偿,也是佳能自愿选择的“体面离开”。(新京报) 佳能撤离珠海,不宜说成“恶意赔偿” 佳能关闭珠海工厂,提供优厚补偿,引发争议。 1 2 3 2
@@
Sentiment Analysis (FinSA) Determine the mood of the sentences in the financial news article provided, identifying whether the mood in them is positive, negative, or neutral. Simply answer ’positive’, ’neutral’, or ’negative’. Up space is exceptionally huge, China Railway assembly, is really the future of the big bull stock ah positive positive 3 3 3 3
@@
Sentiment Matching (FinSM) You need to decide whether the two financial texts given express similar meanings, answer yes or no. Why can’t I pay/ I can’t pay by scanning. yes no 0 1 1 0.7
@@
News Classification (FinNC) Based on the content of the financial report, please determine whether its scope is ’Chinese’, ’international’ or ’foreign’? Bernstein: first Outperform rating on ConocoPhillips (COP.N), \$98 price target. foreign international 1 3 1 1.7
@@
Negative Judgment (FinNJ) Based on the given financial news and entity, you need to determine whether the given entity contains negative news or not, you simply need to answer ’yes’ or ’no’. Since the fall of Tang Xiaosheng, which claimed a transaction volume of 75 billion … Almost every day, there are several online lending companies announced liquidation, cash withdrawal difficulties, the investigation intervention, etc. Entity: Tang Xiaosheng; Domi yes yes 3 3 3 3
@@
Answer Selection (FinAS) Based on the financial question posed, choose the most appropriate of the following four options. Your output should be: ’A’, ’B’, ’C’ or ’D’. If a competitive market is in long-run equilibrium, then all firms () A, use exactly the same production process; B, have the same minimum average cost; C, are economically profitable; D, all of the above. D D 3 3 2 2.7
@@
Stock Prediction (FinSP) Determine whether the company’s stock movement trend is outperforming, neutral, or underperforming based on the impact of the news on the stock data. Gaoxin announced that the real controller intends to reduce its holdings by no more than 6%. Date Opening price Closing price : 2020-09-23, 5.56, 5.52 ; 2020-09-24, 5.47, 5.22; … ; 2020-09-29, 5.02, 5.03 underperformance unpredictable 0 1 2 1
@@
Headline Classification (FinHC) Does the content given below refer to an upward trend in gold commodities? Please answer yes or no. Gold prices rise in India’s major cities yes yes 3 3 2 2.7
@@
Question Answering (FinQA) You will need to analyze financial texts and answer relevant questions based on the content. [YPC] October 21, 2011 - Lodgingsen closed its Series B financing… What are the types of events that occur in the following article? corporate financing indeterminate 0 0 0 0
@@
Event Detection (FinED) Read the following announcements in the financial sector and determine all event types and their corresponding subjects. Please respond in the format :’Event Type, Event Subject’. Where the event type should be in [’credit approval violation’, …’ Suspected Fraud’]. Jinjia shares (002191) due to misrepresentation by the Securities and Futures Commission penalized haoyun science and technology (300448) shareholders disclosure of shareholding reduction plan Letter approval violation, Jinjia shares Letter approval violation, Jinjia shares 3 3 3 3
@@
Entity Recognition (FinER) In the financial statement given below, identify the specific named entity that specifies a person (’PER’), organization (’ORG’), or location (’LOC’).。 During the reporting period, the Company did not have any other material guarantee matters requiring disclosure, except for the financial guarantee business within the scope of operation approved by the China Banking and Insurance Regulatory Commission. China Banking and Insurance Regulatory Commission, ORG China Banking Regulatory Commission, ORG; China, LOC 2 2 3 2.3
@@
Relationship Extraction (FinRE) Carefully analyze the given financial reports and entity pairs, and then among [’Merger’, ’Competition’, …] Choose the option that accurately describes the relationship of the entity pair. Chen Guang, the matchmaker of Zhangyu’s partnership with Custer Group, intends to quit because he thinks the dividend is too small. Entity Pair: Custer Group - Zhang Yu collaborate Impaired partnership 1 2 3 2
@@
Text Summarization (FinTS) Your task is to make a one-sentence summary of the given financial text, keeping it as short as possible. Some time ago, Canon announced the closure of its Zhuhai plant, which had been in operation for 32 years, triggering concern on the Internet…. The generous compensation is also Canon’s voluntary choice to "leave with dignity". (New Beijing News) Canon’s withdrawal from Zhuhai should not be labeled as "malicious compensation". Canon closes Zhuhai plant, offers generous compensation, sparks controversy. 1 2 3 2